Glenn can paint a room in five hours . His father can do it in four hours. How long will it take if they work together
Please help me solve this thankyouuuu

Answer by macston(5194)

.
Glen: 1 room/5 hrs=1/5 room/hr
Father: 1 room/4 hrs=1/4 room/hr
Together: 1/5 room/hr+1/4 room/hr=4/20 room/hr+5/20 room/hr=9/20 room/hr
.
To paint 1 room together at 9/20 room/hr:
1 room/(9/20 room/hr)=2.22 hours
ANSWER: Together they can paint 1 room in 2.22 hours.